OpenStudy (anonymous):

5. The coach of the Morton Valley Softball Team has 6 good hitters and 8 poor hitters. He chooses 3 hitters at random. (a) In how many ways can he choose 2 good hitters and 1 poor hitter? (b) In how many ways can he choose 3 good hitters? (c) In how many ways can he choose at least 2 good hitters?

OpenStudy (anonymous):

hi, a) 120 b) 20 c) 140

OpenStudy (anonymous):

\[\frac{ 6! }{ 2!*4! } *\frac{ 8! }{ 7!*1! } = 120\]

OpenStudy (anonymous):

\[\frac{ 6! }{ 3!*3! } = 20\] this solution for b)